OpenStudy (yb1996):

(sqrt(1 + tan x) - sqrt(1 +sin x))/x^3

OpenStudy (yb1996):

Please find the derivative as x approaches zero

OpenStudy (anonymous):

Do you mean find the limit as \(x\to0\), or find the derivative and evaluate it at 0?

OpenStudy (yb1996):

The first one

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\sqrt{1+\tan x}-\sqrt{1+\sin x}}{x^3}\\ \lim_{x\to0}\frac{\sqrt{1+\tan x}-\sqrt{1+\sin x}}{x^3}\cdot\frac{\sqrt{1+\tan x}+\sqrt{1+\sin x}}{\sqrt{1+\tan x}+\sqrt{1+\sin x}}\\ \lim_{x\to0}\frac{(1+\tan x)-(1+\sin x)}{x^3\left(\sqrt{1+\tan x}+\sqrt{1+\sin x}\right)}\\ \lim_{x\to0}\frac{\tan x-\sin x}{x^3\left(\sqrt{1+\tan x}+\sqrt{1+\sin x}\right)}\\ \lim_{x\to0}\frac{\sin x\left(\dfrac{1}{\cos x}-1\right)}{x^3\left(\sqrt{1+\tan x}+\sqrt{1+\sin x}\right)}\\ \lim_{x\to0}\frac{\sin x}{x}\cdot\lim_{x\to0}\frac{\dfrac{1}{\cos x}-1}{x^2 \left(\sqrt{1+\tan x}+\sqrt{1+\sin x}\right)}\] Here's a preliminary step. Not sure yet if it's in the right direction.

OpenStudy (yb1996):

@SithsAndGiggles Thanks, that really helped.